OpenStudy (anonymous):

On a soccer team, 11 out of 17 players surveyed say they had two or more siblings. The league has 850 players. Which is the best prediction of the number of players in the league that have two or more siblings? (it wont let copy the answers...if u know plzz help.)

OpenStudy (anonymous):

can you plz write the answers? cause this is a pretty hard question about soccer and i know a lot about soccer

OpenStudy (anonymous):

does the question also say how many teams are there in the league?

OpenStudy (anonymous):

lool fine! i'll try.. x/11 = 17/850 x/17 = 11/850 x/850 = 17/11 x/850 = 11/17

OpenStudy (anonymous):

oh ok. this will be easy. just give me a secong

OpenStudy (anonymous):

second

OpenStudy (anonymous):

ohh and I don't think so....hmm..

OpenStudy (anonymous):

i'd say the answer is... D

OpenStudy (anonymous):

did you get it correct?

OpenStudy (anonymous):

okie dokie lets see...

OpenStudy (anonymous):

OMG !!!! YESS!!!

OpenStudy (anonymous):

Your Welcome :D

OpenStudy (anonymous):

OH THANKS @coolnidal

OpenStudy (anonymous):

Anytime :)